OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $14,000. Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $2,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 2% simple annual interest and no down payment. (Hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time. Add the amount of interest to the price of the car.) Which monthly payment amount is lower? Please explain

OpenStudy (shane_b):

For the first plan your payment would be: \[\frac{$14,000-$2,000}{60months}=?\] For the second plan first calculate the interest: \[I=Prt=($14,000)(0.02)(6yrs)=?\]Now add that into the total amount and divide by the number of monthly payments:\[\frac{$14000+I}{72months}=?\]